OpenStudy (hatahateo):

I'll fan and medal Find the real numbers x and y that make the equation true. -4 + yi = x + 3i

OpenStudy (hatahateo):

I already have that -4=x and y=3

OpenStudy (mww):

you need to compare Real and imaginary parts Two complex numbers are equal if and only if the real parts of both match, and the imaginary parts of both also match. so you are correct.

OpenStudy (michele_laino):

hint: when we have an equation between 2 complex numbers: \[a + ib = c + id\] then we can write: \[\left\{ \begin{gathered} a = c \hfill \\ b = d \hfill \\ \end{gathered} \right.\]

OpenStudy (hatahateo):

Is that all I needed to do though? I thought there was one more step

OpenStudy (mww):

nup that's all, complex numbers is generally quite intuitive with operations.

OpenStudy (hatahateo):

Thanks!

OpenStudy (mww):

They do give you longer expressions of this type but the deal is Real(x) = Real(y) and Im(x) = Im(y) is x and y are two equal complex numbers